OpenStudy (anonymous):

Essay: Show all work. A school designer wants to create a whiteboard with the optimal dimensions to enhance learning. It is determined that if one side of the whiteboard is x + y inches, the other side should be inches. Write an algebraic expression for the area of such a whiteboard, simplify it, and include correct units with your solution.

OpenStudy (anonymous):

Seems to me you are missing part of the question. Is this all of it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Essay: Show all work. A school designer wants to create a whiteboard with the optimal dimensions to enhance learning. It is determined that if one side of the whiteboard is x + 2y inches, the other side should be x^2 -2xy+9y^2 inches. Write an algebraic expression for the area of such a whiteboard, simplify it, and include correct units with your solution

OpenStudy (anonymous):

Looks like you are missing that second equation...

OpenStudy (anonymous):

Essay: Show all work. A school designer wants to create a whiteboard with the optimal dimensions to enhance learning. It is determined that if one side of the whiteboard is x + y inches, the other side should be x^2+6xy-4^2 inches. Write an algebraic expression for the area of such a whiteboard, simplify it, and include correct units with your solution.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

no worries.... Area= Side1 times Side2. Side1=x+y. Side 2= x^2+6xy-4^2. Area=(x+y)(x^2+6xy-4y^2) Soooo Area= (x^3+6x^2y-2y^2x)+(x^2y+6xy^2-4y^3) and add like terms to get Area=x^3 + 7 x^2 y + 4 x y^2 - 4 y^3

OpenStudy (anonymous):

is that the whole thing

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

can you plz help with this other one I have

OpenStudy (anonymous):

Essay: Show all work. A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x – 5y ft. and the middle length side x + 6y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out and express in simplest correct mathematical form, including units.